OpenStudy (destinykiara99):

Write a function in simplified form with the following zeros x=-4, 2+i, 2-i

OpenStudy (mightychondrion):

Try doing the conjugate roots first, then multiply the resultant quadratic by (x-4)

OpenStudy (mightychondrion):

Recall that a quadratic can be written as \[ax ^{2}+bx+c\], where the coefficient of b is the sum of the quadratics roots, and the coefficient of c is the product of the quadratics roots, that should help for multiplying the complex roots, then the last real root should be easy. It should look something like \[ax ^{2}\] + (sum of roots [(2+i)+(2-i)])x + (product of roots [(2+i)*(2-i)]).

OpenStudy (destinykiara99):

@mightychondrion that doesn't make sense to me, the whole topic of this question is bery confusing to me

OpenStudy (destinykiara99):

@hartnn could you help?

hartnn (hartnn):

if 'a' is a zero of a polynomial, then, (x-a) is a factor of that polynomial. does this make sense?

OpenStudy (destinykiara99):

Yeah, that makes sense.

hartnn (hartnn):

ok, so if x =-4 is a zero, what is the factor?

OpenStudy (destinykiara99):

(x-(-x)) so (x+4) ?

OpenStudy (destinykiara99):

-4*

hartnn (hartnn):

(x-(-4)) = (x+4)

hartnn (hartnn):

how about if 2+i is a zero?

OpenStudy (destinykiara99):

(x-2+i) ?

hartnn (hartnn):

don't forget the brackets!

hartnn (hartnn):

\((x-(2+i) = (x-2-i)\) ok?

hartnn (hartnn):

how about if x=2-i is a zero?

OpenStudy (destinykiara99):

yeah, I got it. (x-(2-i)) = (x-2-i)

hartnn (hartnn):

(x-(2-i)) is actually (x-2+i)

OpenStudy (destinykiara99):

I realized that after I posted it

hartnn (hartnn):

ok, so you now have the 3 factors, just multiply them together!

hartnn (hartnn):

\(f(x) = (x+4)(x-2-i)(x-2+i) = ...\)

OpenStudy (destinykiara99):

@hartnn I would just combine like terms to simplify that and it would be the final answer correct?

hartnn (hartnn):

yes, thats correct! i suggest you start with \((x-2-i)(x-2+i)\) first :)

OpenStudy (destinykiara99):

simplified would be x^2+4+i?

hartnn (hartnn):

do you know what \((a+b)(a-b)\) equal?

OpenStudy (destinykiara99):

No, what would be A and B in this equation?

hartnn (hartnn):

a= x-2 b = i

hartnn (hartnn):

\((a+b)(a-b) = a^2-b^2\)

OpenStudy (destinykiara99):

x^2+4-i^2 ?

OpenStudy (destinykiara99):

@hartnn